OpenStudy (anonymous):

what is the equation that is perpendicular to y=-3/2x+4 AND that passes through the points (3,9) and how do you get it?

OpenStudy (jdoe0001):

so...what do you think is the slope of => y=-3/2x+4 ?

OpenStudy (anonymous):

-1.5

OpenStudy (jdoe0001):

well... -3/2 yes or -1.5 so a line perpendicular to that one, will have a NEGATIVE RECIPROCAL slope to that one meaning \(\bf -\cfrac{3}{2}\qquad negative\to +\cfrac{3}{2}\qquad reciprocal\to +\cfrac{2}{3}\) so you're really being asked to find the equation of a line whose slope is \(\frac{2}{3}\) and that passes through (3,9) so \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ 9}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}=\cfrac{2}{3} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

oh okay! thanks that helps